OpenStudy (anonymous):

simplify (1+sin(-x)/1+cos(-x))-(1-sin(-x)/1-cos(-x))

OpenStudy (anonymous):

\[\frac{ 1+\sin(-x) }{ 1+\cos(-x) }-\frac{ 1-\sin(-x) }{ 1-\cos(-x) }\]

OpenStudy (anonymous):

Sine and cosine are odd/even functions, respectively, which means \[\sin(-x)=-\sin x\\ \cos(-x)=\cos x\] So, \[\frac{ 1+\sin(-x) }{ 1+\cos(-x) }-\frac{ 1-\sin(-x) }{ 1-\cos(-x) }=\frac{ 1-\sin x }{ 1+\cos x }-\frac{ 1+\sin x }{ 1-\cos x }\]